# An explanation for the absence of echoes in black hole light curve autocorrelations

### v1

@@ Abstract
In the meantime, one can already access the black hole light curves observed over many frequencies and for many sources, including active galactic nuclei, X-ray binaries, and gamma-ray bursts Ulrich et al. (1997); Remillard and McClintock (2006); Berger (2014). Such light curves may be regarded as (single-pixel) “images” that are completely spatially averaged (over both radius and angle). Based on the preceding discussion, one would expect the temporal autocorrelation of many of these light curves to display multiple peaks, with each successive one demagnified by $e^{-\gamma}$ and delayed in time by the characteristic interval $\tau$ between light echoes Fukumura and Kazanas (2008); Moriyama et al. (2019); Wong (2021); Hadar et al. (2021, 2023). However, such time autocorrelations have never been detected.
@@
Where are the light echoes?
In this letter, we revisit the theoretical expectations for the time autocorrelation of a black hole light curve and argue that the peaks caused by lensed images of the main emission are only present if the characteristic timescale $\lambda_{0}$ of temporal correlations in the source is much shorter than the light echo time delay $\tau$. By contrast, if $\lambda_{0}\gtrsim\tau$, then these peaks ought to be absent, even if the lensed images are present and contribute flux to the light curve.
@@
We validate our model against simulated light curves of a stochastic accretion model that we ray trace using a general-relativistic code, and then we fit the model to an observed light curve for Sgr A. We infer that $\lambda_{0}>\tau$, thereby providing an explanation for the absence of light echoes in the time autocorrelations of Sgr A light curves.
@@
As reported in the Supplemental Material (SM), we observe these two behaviors in our numerical simulations. In the intermediate regime $\tau\sim\lambda_{0}$, we find that as $\tau$ increases, the secondary peak changes from a “bump” to an “excess” and usually disappears well before $\tau\approx\lambda_{0}$.
@@
This semi-analytical approach to producing black hole movies with an analytically known covariance function is arguably the “optimal setup” for extracting echoes from light curves, since this method gives us complete control over all correlation scales and black hole parameters. We refer readers to the SM for details of the implementation.
@@
We consider a Kerr black hole with spin $a_{*}=94\%$ observed from an inclination $\theta_{\rm o}=20^{\circ}$, and sources with different characteristic timescales $\lambda_{0}$. We compare their resulting autocorrelations $\mathscr{C}(\Delta t)$ to the prediction (11). The blue line in Fig. 1 (top panel) shows the light curve (6) corresponding to a Keplerian flow with $\lambda_{0}=2\pi/\Omega_{\rm K}$, i.e., proportional to $r_{\rm s}^{3/2}$, with $r_{\rm s}$ the equatorial radius in the source. For this position-dependent correlation time, the resulting GRF is inhomogeneous and anisotropic (see the SM for an example snapshot). The blue open crosses in Fig. 1 (bottom-left panel) plot the corresponding time autocorrelation, which does not present a secondary peak at $\Delta t\sim\tau$, consistent with Eq. (13) and its implications.
@@
When the source has constant correlation scales, the light curve $\mathscr{L}_{0}(t_{\rm o})$ of the direct image $I_{0}(t_{\rm o},\alpha,\beta)$ is a Matérn field with $d=1$, $\nu=3/2$ and correlation length $\lambda_{0}$ (see the SM for a derivation). Its autocorrelation is
@@
For (ii), we compute the autocorrelation of an observed intensity light curve for Sgr A and call upon the intuition built from our model to interpret the results. Specifically, we use the April 7th, 2017 data from the $229.1$ GHz (HI) ALMA (A1) band Wielgus et al. (2022). In Fig. 1 (top panel), we show the observed Sgr A light curve (orange points), as well as a simulated light curve computed from an M87-like GRMHD simulation (green triangles) of a magnetically arrested disk (MAD) with $r_{\rm high}=40$ around a black hole with $a_{*}=85\%$ and $\theta_{\rm o}=163^{\circ}$ Wong et al. (2024); see Refs. Wong et al. (2022, 2024) for more details. To obtain an autocorrelation from the observed Sgr A light curve, we must account for its gaps. As in Ref. Wielgus et al. (2022), we compute a locally normalized discrete correlation function (LNDCF) Lehar et al. (1992); Edelson and Krolik (1988) and use Gaussian process regression (GPR) Rasmussen and Williams (2006) to interpolate the data. This then allows us to apply the same procedure used for the synthetic light curves. The SM provides details of these implementations. The resulting autocorrelations are similar regardless of the method used to obtain them.
@@
Consistent with Ref. Wielgus et al. (2022), Fig. 1 (bottom-right panel) shows no clear signs of a correlation peak at lag $\Delta t\sim\tau$ that could be interpreted as an effect of lensing by Sgr A. We plot the complete autocorrelation in the SM. Since $\mathscr{C}(\tau)\approx 1$, the plot strongly suggests that $\lambda^{\rm{Sgr\,A^{*}}}_{0}>\tau$: that is, the characteristic timescale of correlations in the plasma around Sgr A appears to exceed the time delay between light echoes. Hence, as explained below Eq. (13), we should not expect to see secondary correlation peaks, explaining their absence from observations.
@@
Discussion—We have developed a simple analytical model for the autocorrelation of black hole light curves that offers insight into some of the challenges involved in separating the effects of the plasma from those of the spacetime geometry. When applied to an observed light curve of Sgr A, our model suggests that the temporal correlations inherent in its surrounding plasma suppress the light echoes that are expected from lensing around the black hole, explaining the absence of such signatures.
These results indicate that the inference of black hole parameters from strong lensing effects will be difficult using light curve measurements alone, and likely require future space-VLBI observations that spatially resolve the photon ring. Planning for such observations is underway.
We thank Maciek Wielgus and George Wong for their valuable comments, and for providing the data for the Sgr A and GRMHD light curves, respectively. We also thank Suvendu Giri, Lennox Keeble, Aviad Levis, Leo Stein and Sam Gralla for helpful discussions. A.C.-A. acknowledges support from the Simons Foundation. AL was supported in part by NSF grant 2307888. Some of the simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University.
@@ Appendix A Variation of the photon ring critical parameters with black hole spin and inclination
As discussed in the main text, the presence of a black hole in the midst of an astrophysical source such as a hot radiating plasma is expected to produce light echoes in its observational appearance. [^2] The light echoes that we study herein are produced by orbiting light that is lensed into the photon ring. Such echoes are different from those produced by X-ray “reverberations” that are reflected from different parts of the accretion disk before traveling directly to the observer (along $n=0$ rays), and for which there is now empirical evidence Wilkins et al. (2021). The time delay between such echoes is not controlled by the period of photon orbits, but rather by the distance between the parts of the disk reflecting X-ray emission. The time interval between light echoes can depend on the details of the source, as well as on the black hole mass, spin, and inclination. However, for strongly lensed echoes, the specific value of this time delay is fairly insensitive to these parameters. For instance, as shown in the left panel of Fig. 2, the time lapse $\tau$ incurred per half orbit is around $16M$ for most values of spin and inclination Fukumura and Kazanas (2008); Moriyama et al. (2019); Gralla and Lupsasca (2020), only exceeding $30M$ when the black hole is near extremality and observed from high inclination.
@@ C.2 Position-dependent correlations
As mentioned in the main text, when these parameters are not constant, we cannot derive an analytical expression for the autocorrelation such as the one given in Eq. (14). We can, however, fit the model in Eq. (14) to the numerically computed autocorrelation to obtain an effective timescale $\lambda_{0}$ of temporal correlations. For $\lambda_{0}=2\pi/\Omega_{\rm K}$, $\lambda_{1}=5r_{\rm s}$, and $\lambda_{1}=0.5r_{\rm s}$ (a snapshot of the resulting field is shown in the right panel of Fig. 3), we find an effective value for the correlation timescale of $\lambda_{0}=88M$, as shown in Fig. 4 with the blue lines and points. This effective timescale of the source correlations is expected to be different from the one inferred from observations of the source lensed by the black hole ($\lambda_{0}^{\rm fitted}=54M$), whose presence relativistically broadens the autocorrelation. Although this broadening does not provide a clear signature of the light echo (nor a correlation at the relevant timescale), if one knew the time correlation of the underlying emission, then one could indirectly infer the presence of the black hole due to this shift.

### v2

# Explanation for the absence of secondary peaks in black hole light curve autocorrelations

@@ Abstract
In the meantime, one can already access the black hole light curves observed over many frequencies and for many sources, including active galactic nuclei, X-ray binaries, and gamma-ray bursts Ulrich et al. (1997); Remillard and McClintock (2006); Berger (2014). Such light curves may be regarded as (single-pixel) “images” that are completely spatially averaged (over both radius and angle). Based on the preceding discussion, one would expect the temporal autocorrelation of many of these light curves to display multiple peaks, with each successive one demagnified by $e^{-\gamma}$ and delayed in time by the characteristic interval $\tau$ between light echoes Fukumura and Kazanas (2008); Moriyama et al. (2019); Chesler et al. (2021); Wong (2021); Hadar et al. (2021, 2023). However, such time autocorrelations have never been detected.
@@
Where are the light echo autocorrelation peaks?
In this letter, we revisit the theoretical expectations for the time autocorrelation of a black hole light curve and argue that the peaks caused by lensed images of the main emission are only present if the characteristic timescale $\lambda_{0}$ of temporal correlations in the source is much shorter than the light echo time delay $\tau$. By contrast, if $\lambda_{0}\gtrsim\tau$, then these maxima ought to be absent, even if the lensed images are present and contribute flux to the light curve.
@@
We validate our model against simulated light curves of a stochastic accretion model that we ray trace using a general-relativistic code, and then we fit the model to an observed light curve for Sgr A. We infer that $\lambda_{0}>\tau$, providing an explanation for the absence of secondary peaks in the time autocorrelations of Sgr A light curves.
@@
As reported in the Supplemental Material (SM) SM , we observe these two behaviors in our numerical simulations. In the intermediate regime $\tau\sim\lambda_{0}$, we find that as $\tau$ increases, the secondary peak changes from a “bump” to an “excess” and usually disappears well before $\tau\approx\lambda_{0}$.
@@
This semi-analytical approach to producing black hole movies with an analytically known covariance function is arguably the “optimal setup” for extracting echoes from light curves, since this method gives us complete control over all correlation scales and black hole parameters. See the SM for the details of the implementation SM .
@@
We consider a Kerr black hole with spin $a_{*}=94\%$ observed from an inclination $\theta_{\rm o}=20^{\circ}$, and sources with different characteristic timescales $\lambda_{0}$. We compare their resulting autocorrelations $\mathscr{C}(\Delta t)$ to the prediction (11). The blue line in Fig. 1 (top panel) shows the light curve (6) corresponding to a Keplerian flow with $\lambda_{0}=2\pi/\Omega_{\rm K}$, i.e., proportional to $r_{\rm s}^{3/2}$, with $r_{\rm s}$ the equatorial radius in the source. For this position-dependent correlation time, the resulting GRF is inhomogeneous and anisotropic (see SM for an example snapshot SM ). The blue open crosses in Fig. 1 (bottom-left panel) plot the corresponding time autocorrelation, which does not present a secondary peak at $\Delta t\sim\tau$, consistent with Eq. (13) and its implications.
@@
If the source has constant correlation scales, then the light curve $\mathscr{L}_{0}(t_{\rm o})$ of the direct image $I_{0}(t_{\rm o},\alpha,\beta)$ is a Matérn field with $d=1$, $\nu=3/2$ and correlation length $\lambda_{0}$ SM . Its autocorrelation is
@@
For (ii), we compute the autocorrelation of an observed intensity light curve for Sgr A and call upon the intuition built from our model to interpret the results. Specifically, we use the April 7th, 2017 data from the $229.1$ GHz (HI) ALMA (A1) band Wielgus et al. (2022). In Fig. 1 (top panel), we show the observed Sgr A light curve (orange points), as well as a simulated light curve computed from an M87-like GRMHD simulation (green triangles) of a magnetically arrested disk (MAD) with $r_{\rm high}=40$ around a black hole with $a_{*}=85\%$ and $\theta_{\rm o}=163^{\circ}$ Wong et al. (2024); see Refs. Wong et al. (2022, 2024) for more details. To obtain an autocorrelation from the observed Sgr A light curve, we must account for its gaps. As in Ref. Wielgus et al. (2022), we compute a locally normalized discrete correlation function (LNDCF) Lehar et al. (1992); Edelson and Krolik (1988) and use Gaussian process regression (GPR) Rasmussen and Williams (2006) to interpolate the data. This then allows us to apply the same procedure used for the synthetic light curves. See the SM for details of these implementations SM . The resulting autocorrelations are similar regardless of the method used to obtain them.
@@
Consistent with Ref. Wielgus et al. (2022), Fig. 1 (bottom-right panel) shows no clear signs of a correlation peak at lag $\Delta t\sim\tau$ that could be interpreted as an effect of lensing by Sgr A.
We plot the complete autocorrelation in the SM SM . Since $\mathscr{C}(\tau)\approx 1$, the plot strongly suggests that $\lambda^{\rm{Sgr\,A^{*}}}_{0}>\tau$: that is, the characteristic timescale of correlations in the plasma around Sgr A appears to exceed the time delay between light echoes. Hence, in accordance with Eq. (13), we should not expect to see secondary correlation peaks, explaining their absence from observations.
@@
Discussion—We have developed a simple analytical model for the autocorrelation of black hole light curves that offers insight into some of the challenges involved in separating the effects of the plasma from those of the spacetime geometry. When applied to an observed light curve of Sgr A, our model suggests that the temporal correlations inherent in its surrounding plasma suppress the autocorrelation peaks expected from lensing around the black hole, explaining the absence of such signatures.
These results indicate that the inference of black hole parameters from strong lensing effects will be difficult via light curve autocorrelations alone, and likely require future space-VLBI observations that spatially resolve the photon ring. Planning for such observations is underway.
We thank Maciek Wielgus and George Wong for their valuable comments, and for providing the data for the Sgr A and GRMHD light curves, respectively. We also thank Neal Dalal, Suvendu Giri, Lennox Keeble, Aviad Levis, Leo Stein and Sam Gralla for helpful discussions. CG and AL also thank the Aspen Center for Physics, which is supported by NSF grant PHY-2210452. ACA acknowledges support from the Simons Foundation. AL is supported by NSF grants AST-2307888 and PHY-2340457. CG was supported in part by the IBM Einstein Fellow Fund at the Institute for Advanced Study. Some of the simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University.
@@ Appendix A Variation of the photon ring critical parameters with black hole spin and inclination
As discussed in the main text, the presence of a black hole in the midst of an astrophysical source such as a hot radiating plasma is expected to produce light echoes in its observational appearance. [^2] The light echoes that we study herein are produced by orbiting light that is lensed into the photon ring. Such echoes are different from those produced by X-ray “reverberations” that are reflected from different parts of the accretion disk before traveling directly to the observer (along $n=0$ rays), and for which there is now empirical evidence Wilkins et al. (2021). The time delay between such echoes is not controlled by the period of photon orbits, but rather by the distance between the parts of the disk reflecting X-ray emission. The time interval between light echoes can depend on the details of the source, as well as on the black hole mass, spin, and inclination. However, for strongly lensed echoes, the specific value of this time delay is fairly insensitive to these parameters. For instance, as shown in the left panel of Fig. 2, the time lapse $\tau$ incurred per half orbit is around $16M$ for most values of spin and inclination Fukumura and Kazanas (2008); Moriyama et al. (2019); Gralla and Lupsasca (2020), only exceeding $30M$ when the black hole is near extremality and observed from high inclination. Assuming that $\tau$ takes a single value is a conservative approach. In practice, different sources produce a distribution of time delays in low-order images, broadening out the time autocorrelation and thereby obscuring light echo peaks.
@@ C.2 Position-dependent correlations
As mentioned in the main text, when these parameters are not constant, we cannot derive an analytical expression for the autocorrelation such as the one given in Eq. (14). We can, however, fit the model in Eq. (14) to the numerically computed autocorrelation to obtain an effective timescale $\lambda_{0}$ of temporal correlations. For $\lambda_{0}=2\pi/\Omega_{\rm K}$, $\lambda_{1}=5r_{\rm s}$, and $\lambda_{1}=0.5r_{\rm s}$ (a snapshot of the resulting field is shown in the right panel of Fig. 3), we find an effective value for the correlation timescale of $\lambda_{0}=88M$, as shown in Fig. 4 with the blue lines and points. This effective timescale of the source correlations is expected to be different from the one inferred from observations of the source lensed by the black hole ($\lambda_{0}^{\rm fitted}=54M$), whose presence relativistically broadens the autocorrelation. Although this broadening does not provide a clear signature of the light echo (nor a correlation at the relevant timescale), if one knew the time correlation of the underlying emission, then one could indirectly infer the presence of the black hole due to this shift; for further discussion of this point, see Sec. E below.
@@ Appendix D Autocorrelations from observations of Sagittarius A*
Appendix E Power spectrum of light curve autocorrelations
Here, we briefly discuss the characteristic signatures that light echoes introduce in the power spectrum (Fourier transform) of the time autocorrelation $\mathcal{C}(\Delta t)$. Suppose as in Eq. (11) that the observed light curve has autocorrelation
$\mathscr{C}(\Delta t)=\mathscr{C}_{0}(\Delta t)\mathopen{}\mathclose{{}\left[1-2e^{-\gamma}\mathscr{C}_{0}(\tau)}\right]+e^{-\gamma}\mathopen{}\mathclose{{}\left[\mathscr{C}_{0}(\Delta t+\tau)+\mathscr{C}_{0}(\Delta t-\tau)}\right]+\mathcal{O}\mathopen{}\mathclose{{}\left(e^{-2\gamma}}\right),$ (32)
where $\mathscr{C}_{0}(\Delta t)$ denotes the time autocorrelation of the direct ($n=0$) emission, with Fourier transform
$\tilde{\mathscr{C}}_{0}(k)=\int\mathscr{C}_{0}(\Delta t)e^{-ik\Delta t}\mathop{}\!\mathrm{d}\Delta t,\qquad\mathscr{C}_{0}(\Delta t)=\frac{1}{2\pi}\int\tilde{\mathscr{C}}_{0}(k)e^{+ik\Delta t}\mathop{}\!\mathrm{d}k.$ (33)
Then, the Fourier transform of the full autocorrelation $\mathscr{C}(\Delta t)$ is
$\tilde{\mathscr{C}}(k)$ $=\tilde{\mathscr{C}}_{0}(k)\mathopen{}\mathclose{{}\left[1-2e^{-\gamma}\mathscr{C}_{0}(\tau)}\right]+e^{-\gamma}\mathopen{}\mathclose{{}\left[e^{+ik\tau}\tilde{\mathscr{C}}_{0}(k)+e^{-ik\tau}\tilde{\mathscr{C}}_{0}(k)}\right]+\mathcal{O}\mathopen{}\mathclose{{}\left(e^{-2\gamma}}\right).$ (34)
Thus, we see that the autocorrelation power spectrum in the presence of the black hole, $\tilde{\mathscr{C}}(k)$, differs from the power spectrum in the absence of a black hole, $\tilde{\mathscr{C}}_{0}(k)$, by a multiplicative term
$\frac{\tilde{\mathscr{C}}(k)}{\tilde{\mathscr{C}}_{0}(k)}=1+2e^{-\gamma}\mathopen{}\mathclose{{}\left[\cos\mathopen{}\mathclose{{}\left(\tau k}\right)-\mathscr{C}_{0}(\tau)}\right]+\mathcal{O}\mathopen{}\mathclose{{}\left(e^{-2\gamma}}\right).$ (35)
This calculation only accounts for the first ($n=1$) light echoes. We expect the $n^{\text{th}}$-order light echoes to introduce additional modes oscillating with frequency $n\tau$ but exponentially suppressed by $e^{-n\gamma}$.
These characteristic oscillations in the power spectrum may provide a way to detect the presence of light echoes in black hole light curves even when they do not produce secondary peaks in the time autocorrelation. In practice, however, we expect a detection of the signature (35) to be very challenging for a number of reasons, including:
  • First, as mentioned in Sec. A, different sources can produce a distribution of time delays in low-order (small-$n$) images, especially at nonzero inclination, so the oscillation $\cos\mathopen{}\mathclose{{}\left(\tau k}\right)$ will be smeared over a range of $\tau$.
  • Second, the intrinsic correlation structure of the source cannot be directly observed, so that one can directly access only $\mathscr{C}(\Delta t)$ or $\tilde{\mathscr{C}}(k)$, but not $\mathscr{C}_{0}(\Delta t)$ nor $\tilde{\mathscr{C}}_{0}(k)$. In other words, the quantity in Eq. (35) is unobservable.
  • Third, the source power spectrum $\tilde{\mathscr{C}}_{0}(k)$ may itself have oscillations with frequencies near $\tau$. Intrinsic correlations on the same timescale $\lambda\sim\tau$ as the lensing echoes may very well be present in the accretion flow.
To sum up, one would have to know the details of the astrophysical source, and in particular its correlation structure $\tilde{\mathscr{C}}_{0}(k)$, to claim that an observed oscillation with frequency $\tau$ were produced by lensing rather than the source itself.